OpenStudy (anonymous):

What is the value of a? 5(a-4)-8a=55

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Start by distributing  What do you get?

OpenStudy (anonymous):

???

pooja195 (pooja195):

What do you get once you distribute?

pooja195 (pooja195):

5 x a=? 5 x -4=?

OpenStudy (anonymous):

5 x -4=-20

pooja195 (pooja195):

ok good \[\huge~\rm5a−20−8a=55\] combine like terms is the next step \[\huge~\rm~(5a−8a)+(−20)=55\]

pooja195 (pooja195):

5-8= =?

OpenStudy (anonymous):

-3???????

pooja195 (pooja195):

Yes :) \[\huge~\rm~−3a−20=55 \] From here solve for a im sure you can do it :)

OpenStudy (anonymous):

-3a=55 am i correct so far?

pooja195 (pooja195):

you need to add 20 to both sides...

OpenStudy (anonymous):

hmm i did it wrong

pooja195 (pooja195):

Try again

pooja195 (pooja195):

\[\huge~\rm~−3a−20+20=55+20\]

OpenStudy (anonymous):

-3a-40=75?

pooja195 (pooja195):

-20+20=0 55+20=?

OpenStudy (anonymous):

ohh -20 oooppsss

OpenStudy (anonymous):

-3a=75

pooja195 (pooja195):

Good now divide both sides by -3 -3/-3=1 so just a 75/-3=?

OpenStudy (anonymous):

11.6?

pooja195 (pooja195):

Nope... what is 75 divided by -3?

OpenStudy (anonymous):

wait i used the wrong number

OpenStudy (anonymous):

-25

pooja195 (pooja195):

Well done :)

OpenStudy (anonymous):

YOU ARE SOOOO AWESOME THANK YOU

pooja195 (pooja195):

yw ^_^